If today is my birthday, and my age fifteen years ago is the square of my age fifteen years from now, what is my age?
Answer by ankor@dixie-net.com(22740)


I think it should read:
my age fifteen years ago squared, is my age fifteen years from now,
what is my age?
:
Let a = your present age
(a-15)^2 = (a+15)
a^2 - 30a + 225 = a + 15
a^2 - 30a - a + 225 - 15 = 0
a^2 - 31a + 210 = 0
Factors to
(a-10)(a-21) = 0
Two solutions
a = 10
a = 21, this is the only solution that makes sense
:
:
Check
(21-15)^2 = 36
21 + 15 = 36